### v1

# Revisiting Koehler’s experiment of measuring the ratio of the specific heats of air by self-sustained oscillations: a more concise theoretical interpretation

Abstract
We revisit Koehler’s experiment, a clever modification of Rüchardt’s experiment designed to measure the ratio of specific heats of gas. The theory of self-sustained oscillations in Koehler’s experiment was provided by Koehler (1950). However, its lengthy and dense analysis may pose challenges to readers due to the complexity of the calculations. Following Koehler’s approximation for pressure changes, we explicitly present the model equations as piecewise linear differential systems and qualitatively analyze the periodic solutions from a geometric perspective. This concise and transparent approach addresses a fundamental question about Koehler’s experiment: why is the oscillation frequency nearly equal to the Rüchardt frequency? Our analysis avoids intricate calculations and will be particularly helpful for teachers and students who encounter Koehler’s experiment in general physics laboratory classes.
I Introduction
Koehler’s experiment for measuring the ratio of specific heats $\gamma=C_{p}/C_{v}$ of air (or other gases) is a modification of Rüchardt’s experiment, a classic experiment widely used in general physics laboratory courses Koehler1951 ; Lange2000 . As shown in Fig. 1(a), the apparatus of Rüchardt’s experiment consists of a large vessel with a vertical tube containing a closely fitting steel ball. To facilitate the discussion that follows, it is useful to briefly recall the theory of Rüchardt’s experiment. When the ball is stationary at its equilibrium position, the equilibrium pressure satisfies the equation $P_{0}=B+\frac{mg}{\pi R^{2}}$, Where $B$ is atmospheric pressure, mm and $R$ are the mass and radius of the ball. When the ball is displaced from its equilibrium position and released, its rapid motion induces pressure fluctuations due to the adiabatic compression or expansion of the gas within the container. For the ball, we have
$m\ddot{x}=-\pi R^{2}\cdot\Delta P$ (1)
Where $x$ is the displacement of the ball from its equilibrium position, and $\Delta P$ is the gas pressure change in the vessel. Considering the radiabatic equation $PV^{\gamma}=const.$,
we have
$\Delta P=-\frac{\gamma P\Delta V}{V}\sim-\frac{\gamma P_{0}\Delta V}{V_{0}}=-\frac{\gamma P_{0}\cdot\pi R^{2}}{V_{0}},$ (2)
where $V_{0}$ is the volume of the vessel. Substituting it into Eq.(1), it can be seen that the motion of the ball is simple harmonic oscillation. The relationship between the period of oscillation and $\gamma$ is given by
$\gamma=\frac{4mV_{0}}{T^{2}P_{0}R^{4}}.$ (3)
Although Rüchardt’s method has the advantages of simplicity and mechanical intuitiveness, it is not as simple as it first appears. Unavoidable frictional damping and gas leakage limit the number of oscillations, typically to only a few, preventing precise measurement of the period. In 1951, Koehler described an interesting modification involving a gas-feeding pump and introducing a small hole in the tube (as shown in Fig. 1(b))Koehler1951 . The ball is lifted by gas overpressure and falls back due to the pressure drop from gas leakage through the hole. By carefully adjusting the gas flow, the ball can oscillate symmetrically about the hole indefinitely. The oscillations closely approximate harmonic motion, and the relationship between the period and $\gamma$ is still described by Eq.(3). The advantages of Koehler’s method are evident: it allows prolonged measurement of the period, thereby improving the ease and precision of measurements. Therefore, compared to Rüchardt’s experiment, Koehler’s experiment offers greater advantages for teaching purposes.
Figure 1: (a) and (b) Schematic diagrams of Rüchardt apparatus and Koehler apparatus. (c) An image of experimental apparatus.
In China, many universities incorporate the Koehler experiment into general physics laboratory courses. However, there is limited evidence of its widespread adoption globally. A search of online resources reveals few instances of its use in English-speaking countries, particularly in American colleges and universities. Even in experimental demonstrations of the specific heat ratio of gases on YouTube, nearly all are based on Rüchardt’s design. By incorporating a simple air pump (costing less than \$10), the Koehler experiment achieves sustained oscillation through a clever design, enabling more precise measurement of the gas-specific heat ratio. Thus, we argue that the Koehler experiment provides significant pedagogical advantages over the Rüchardt experiment in student laboratory teaching and has the potential for broader adoption in other regions in the future. Furthermore, even in schools that currently use the Rüchardt experiment, instructors could easily adapt the setup to implement the Koehler version or allow students to explore it as an independent project, given its simplicity. For reference, the detailed parameters of the experimental apparatus employed in this study are provided in Appendix A.
In courses using Koehler’s experiment, the theoretical introduction typically follows the framework of the original Rüchardt experiment. Students are generally informed that the ball’s motion approximates harmonic behavior, as described by Eq.(3). However, an issue arises: the supply and leakage of gas indicate that the pressure fluctuations within the container are not solely determined by the adiabatic changes in gas volume. The pressure change corresponding to the restoring force is no longer linearly proportional to the displacement of the ball. Why do these motions still closely approximate harmonic oscillations and satisfy Eq.(3)? In fact, the ball’s motion represents a nonlinear, self-sustained oscillation. Koehler presented a theoretical explanation for such self-sustained oscillations in another paper Koehler1950 . Assuming the ball’s motion to be sinusoidal, he used nonlinear dynamic equations to demonstrate that the pressure fluctuations are also approximately sinusoidal, with the oscillation frequency closely matching the Rüchardt frequency. However, this analysis is indirect, with a lengthy and dense derivation, potentially posing challenges to readers given the complexity of the calculations. To our knowledge, no other studies provide a more accessible theoretical analysis of this issue. This might explain why textbooks often mention it briefly, stating, for example, that “the pressure changes caused by gas supply and leakage are relatively small, so…” without offering detailed explanations.
The following sections of this paper will revisit this issue and provide a concise theoretical interpretation of the self-sustained oscillations. The paper is organized as follows. In section 2, following Koehler’s approximation for pressure changes, we explicitly present the dynamic equations. In section 3, we dimensionless the model equations and expressed them as 3D piecewise linear differential systems, and analyzed the behavior of the systems near their fixed points in a 2D projection plane. In section 4, we qualitatively analyze the properties of periodic solutions of the system from a geometric perspective, demonstrating that the oscillation frequency in Koehler’s experiment is nearly equal to the Rüchardt frequency.
II Dynamic equations of self-sustained oscillations
Let $x$ and $v$ denote the vertical coordinate and velocity of the ball, $P$ the pressure in the vessel (assumed spatially uniform) which is a function of the ball’s coordinate $x$ and time $t$. The dynamic equations are
$\frac{dx}{dt}=v,$ (4a)
$\frac{dv}{dt}=\frac{(P-P_{0})\pi R^{2}}{m},$ (4b)
$\frac{dP}{dt}=\frac{\partial P}{\partial x}\frac{dx}{dt}+\frac{\partial P}{\partial t}.$ (4c)
Eq.(4c) shows that the rate of pressure change consists of two parts: the term ${\frac{\partial P}{\partial x}}{\frac{dx}{dt}}$ caused by the adiabatic change of gas volume due to the motion of the ball, and the term $\frac{\partial P}{\partial t}$ contributed by gas supply and leakage. For the adiabatic process of an ideal gas,
$\frac{\partial P}{\partial x}=\frac{\partial P}{\partial V}\frac{\partial V}{\partial x}=-\frac{\gamma P\pi R^{2}}{V}\sim-\frac{\gamma P_{0}\pi R^{2}}{V_{0}}.$ (5)
For $\frac{\partial P}{\partial t}$, here we use the approximation from Koehler’s paperKoehler1950 . The contribution to $\frac{\partial P}{\partial t}$ due to gas supply is approximated as a positive constant
$\frac{\partial P}{\partial t}_{supply}=jB.$ (6)
The contribution to $\frac{\partial P}{\partial t}$ due to the gas leakage is approximated as proportional to the difference in pressure between the inside and outside of the container.
$\frac{\partial P}{\partial t}_{leakage}=-g(x)(P-B),\quad g(x)=\begin{cases}e+f,&\quad x\geq 0\\ e,&\quad x<0\end{cases},$ (7)
The coordinate origin is defined at the center of the small hole on the tube. When $x<0$, gas escapes solely through the annular gap between the ball and the tube wall. When $x\geq 0$, gas escapes through both the annular gap around the ball and the small hole. The finite size of the small hole is neglected, resulting in a discontinuity of $g(x)$ at 0. However, this effect is minor and not crucial to the main point. By substituting Eqs. (5), (6), and (7) into Eqs. (4), the final dynamic equations are derived as
$\frac{dx}{dt}=v,$ (8a)
$\frac{dv}{dt}=\frac{(P-P_{0})\pi R^{2}}{m},$ (8b)
$\frac{dP}{dt}=-\frac{\gamma P_{0}\pi R^{2}}{V_{0}}v+jB-g(x)(P-B).$ (8c)
When the term jBg(x)(PB)jB-g(x)(P-B) is removed, the equations reduce to those describing harmonic motion in the ideal Rüchardt experiment. The Rüchardt frequency is given by $\omega^{2}=\frac{\gamma P_{0}\pi^{2}R^{4}}{mV_{0}}$(equivalent to Eq. (3)). The term involving the piecewise function $g(x)$ introduces nonlinearity. As a result, the equations can no longer be solved analytically.
III Preliminary qualitative analysis
We first nondimensionalize Eqs. (8) as follows:
$\frac{dy}{d\tau}=u,$ (9a)
$\frac{du}{d\tau}=Ap,$ (9b)
$\frac{dp}{d\tau}=-\frac{1}{A}u-G(y)p+[J-G(y)(p_{0}-1)],$ (9c)
where $\tau=\omega t,y=\frac{x}{R},u=\frac{v}{R\omega},p=\frac{P-P_{0}}{B},p_{0}=\frac{P_{0}}{B},A=\frac{\pi RB}{m\omega^{2}},J=\frac{j}{\omega}$, and
$G(y)=\frac{g(x)}{\omega}=\begin{cases}G_{+}=(e+f)/\omega,\quad&y\geq 0,\\ G_{-}=e/\omega,\quad&y<0.\end{cases}$
In mathematics, Eqs.(9) can be classified as piecewise linear differential systems (PWLD systems). These systems consist of two linear subsystems separated by the plane defined by $\{y=0\}$, referred to as the separation plane. They can be written in matrix form as
$\dot{\mathbf{x}}=\begin{cases}M_{+}\mathbf{x}+\mathbf{e}_{+}&\mathrm{if}\quad y\geq 0\\ M_{-}\mathbf{x}+\mathbf{e}_{-}&\mathrm{if}\quad y<0\end{cases}$ (10)
with
$\mathbf{x}=\begin{pmatrix}y\\ u\\ p\end{pmatrix},\quad M_{+}=\begin{pmatrix}0&1&0\\ 0&0&A\\ 0&-\frac{1}{A}&-G_{+}\end{pmatrix},\quad M_{-}=\begin{pmatrix}0&1&0\\ 0&0&A\\ 0&-\frac{1}{A}&-G_{-}\end{pmatrix}$
and
$\mathbf{e}_{+}=\begin{pmatrix}0\\ 0\\ J-G_{+}(p_{0}-1)\end{pmatrix},\quad\mathbf{e}_{-}=\begin{pmatrix}0\\ 0\\ J-G_{-}(p_{0}-1)\end{pmatrix}$.
PWLD systems are widely used to model real-world phenomena in fields such as electrical circuits, control systems, impact and friction mechanics, and biological systems. However, analyzing periodic solutions, which often correspond to steady-state responses in general PWLD systems, is notoriously challenging Jeffrey2018 . This difficulty arises because PWLD systems autonomously switch their behavior at specific boundaries, leading to discontinuities and distinct sets of governing equations in different regions of the phase space. Accurately handling transitions across these boundaries often requires solving implicit or transcendental equations, which can be particularly challenging or even infeasible in practice. Despite these challenges, the following sections of this paper will provide an approximate analysis to demonstrate if the periodic solutions exist, specifically, solutions in which the small ball passes through the hole position once upward and once downward during a single period, then the system’s motion frequency is indeed very close to the Rüchardt frequency.
Analyzing fixed points (equilibrium points) yields valuable insights for dynamical systems. In both half-space $\{y\geq 0\}$ and $\{y<0\}$, Eqs.(9) do not exhibit global fixed points. However, note that on the right-hand side of the equations, the variable $y$ appears only within $G(y)$, whose constant value is piecewise determined by $y$. Thus, $y$ can be treated as a parameter, allowing for an analysis of the system in the $(u,p)$ projection plane for each half-space, $\{y\geq 0\}$ and $\{y<0\}$.
$\frac{du}{d\tau}=Ap,$ (11)
$\frac{dp}{d\tau}=-\frac{1}{A}u-Gp+[J-G(p_{0}-1)],$
where $G$ represents $G(y)$, with its values in the two half-space $\{y\geq 0\}$ and $\{y<0\}$ denoted as $G_{+}$ and $G_{-}$, respectively. Let $\frac{du}{d\tau}=0,\frac{dp}{d\tau}=0$, we obtain the fixed point of this subsystem, denoted as $(u^{*},p^{*})$, given by
$u^{*}=A\left[J-G(p_{0}-1)\right],\quad p^{*}=0.$
Next, the local behavior around the fixed point is analyzed. Let $u=u^{*}+\eta(\tau),p=p^{*}+\xi(\tau)$, where $\eta(\tau)$ and $\xi(\tau)$ represent small deviations from the fixed point. Substituting these into Eqs. (11), the system is rewritten as
$\frac{d}{d\tau}\binom{\eta}{\xi}=\begin{bmatrix}0&A\\ -\frac{1}{A}&-G\end{bmatrix}\binom{\eta}{\xi}$. (12)
The Jacobian matrix $\begin{bmatrix}0&A\\ -\frac{1}{A}&-G\end{bmatrix}$ has eigenvalues $\lambda=-\frac{G}{2}\pm i\sqrt{1-\frac{G^{2}}{4}}$. Since $G$ is positive and much smaller than $1$ (as shown in Appendix B, which estimates the experimental values of parameters such as $A$, $J$, and $G$ ), the term $1-\frac{G^{2}}{4}$ is positive. Therefore, the fixed point $(u^{*},p^{*})$ is a stable focus, and any trajectory in the plane will spiral towards this point. The perturbations $\eta(\tau)$ and $\xi(\tau)$ can be expressed as
$\eta(\tau)=C_{1}\exp(-\frac{G}{2}\tau)\cos(\sqrt{1-\frac{G^{2}}{4}}\tau+\Phi),$ (13a)
$\xi(\tau)=\frac{1}{A}\frac{du}{d\tau}=-C_{1}\exp(-\frac{G}{2}\tau)\cos(\sqrt{1-\frac{G^{2}}{4}}\tau+\Phi-\varphi),$ (13b)
where $C_{1}$, $\Phi$ are undetermined coefficients related to initial values, and $tan(\varphi)=\frac{\sqrt{1-\frac{G^{2}}{4}}}{\frac{G}{2}}$. These results show that the perturbations in the ball’s velocity and the gas pressure are both characterized by decaying oscillations with a frequency of $\sqrt{1-\frac{G^{2}}{4}}\sim 1$. This indicates that in Koehler’s experiment, whether the ball is below the small hole (in half space y<0{y<0}) or above it (in half space $y\geq 0$), the changes in the ball’s velocity and the gas pressure remain closely approximated by harmonic motion, with a frequency nearly identical to the Rüchardt frequency $\omega$.
It is important to note that, so far, we can only state that the oscillation frequency of the system within each subspace is close to the Rüchardt frequency. However, we cannot arbitrarily say that the frequency of the overall motion, obtained by combining the trajectories from each subspace, is also close to the Rüchardt frequency. This is because the exact duration for which the system remains in each subspace is not known. For instance, let $T_{0}$ represent the Rüchardt period. If the system spends $\frac{T_{0}}{2}$ in one half-space and $\frac{T_{0}}{3}$ in another half-space, which may be possible, the total period of the motion will be $(\frac{T_{0}}{2}+\frac{T_{0}}{3})$, not $T_{0}$, resulting in a frequency that deviates from the Rüchardt frequency. In the following section, we will present a geometric analysis to demonstrate that, under appropriate experimental conditions, when the system alternates between two half-spaces, the overall motion period approaches $T_{0}$, and the oscillation frequency remains close to the Rüchardt frequency.
IV Discussion on Periodic Solutions of Model Equations
In this section, we will estimate the time the system spends in each subspace from a geometric perspective in phase space and demonstrate that the overall motion period approaches $T_{0}$. We find that the original 3D PWLD system with one separation plane $\{y=0\}$ can be equivalently simplified to a 2D PWLD system with a single separation straight line, which is more analytically tractable. Although this section is somewhat lengthy, it essentially describes how the system’s trajectory revolves around two alternating focal points in a periodic manner.
IV.1 Difficulties in analyzing periodic solutions of the 3D PWLD system
Let us first provide a detailed explanation of the difficulty in rigorously analyzing the periodic solutions of system (10). The Poincaré mapping is a powerful tool for studying periodic solutions in dynamical systems. However, for general PWLD systems, constructing a Poincaré map requires careful treatment of transitions across the switching surface within each subspace, which presents significant challenges. For system (10), we treat the separation plane $\{y=0\}$ as the Poincaré section. As discussed previously, when the system operates in any half-space, the velocity of the ball can be represented as
$u=u^{*}+\eta(\tau)=u^{*}+C_{1}\exp(-\alpha\tau)\cos(\Omega\tau+\Phi),$ (14)
where $\alpha=\frac{G}{2}$, $\Omega=\sqrt{1-\frac{G^{2}}{4}}$. According to Eq. (9a), the expression for the coordinates of the ball can be obtained as
$y(\tau)$ $=y_{0}+\int_{0}^{\tau}u(t)dt$ (15)
$=y_{0}+u^{*}\tau+\frac{C_{1}}{\alpha^{2}+\Omega^{2}}\left\{-\alpha e^{-\alpha\tau}\cos(\Omega\tau+\Phi)+\Omega e^{-\alpha\tau}\sin(\Omega\tau+\Phi)+\alpha\cos(\Phi)-\Omega\sin(\Phi)\right\}.$
By solving $y(\tau)=0$, in principle, the state of the system when switching between subspaces and its running time in each subspace can be obtained. However, $y(\tau)=0$ is too complex to yield analytical solutions, which complicates the construction of Poincaré mappings to analyze the periodic solution behavior of the system.
IV.2 Simplified problem: 2D PWLD system with one separation straight line
From the analysis above, it is evident that the trajectory of the system in any half-space, when projected onto the $(u,p)$ plane, appears as a simple spiral around the focus. However, the switching between subspaces is controlled by the condition $y=0$ in the third dimension, which presents significant challenges in solving for the switching state. To avoid this difficulty, we simplify system (10) by reducing its dimensionality, resulting in a purely 2D PWLD system as follows:
$\dot{\mathbf{x}}=\begin{cases}M_{+}\mathbf{x}+\mathbf{e}_{+}&\mathrm{if}\quad p\leq 0\\ M_{-}\mathbf{x}+\mathbf{e}_{-}&\mathrm{if}\quad p>0\end{cases}$ (16)
with
$\mathbf{x}=\begin{pmatrix}u\\ p\end{pmatrix},\quad M_{+}=\begin{pmatrix}0&A\\ -\frac{1}{A}&-G_{+}\end{pmatrix},\quad M_{-}=\begin{pmatrix}0&A\\ -\frac{1}{A}&-G_{-}\end{pmatrix}$,
and
$\mathbf{e}_{+}=\begin{pmatrix}0\\ u^{*}_{+}\end{pmatrix},\quad\mathbf{e}_{-}=\begin{pmatrix}0\\ u^{*}_{-}\end{pmatrix}$,
where
$u^{*}_{+}=A\left[J-G_{+}(p_{0}-1)\right]<0$ (the ball’s motion tends towards the steady state of falling),
$u^{*}_{-}=A\left[J-G_{-}(p_{0}-1)\right]>0$ (the ball’s motion tends towards the steady state of rising).
Figure 2: (a) and (b) The phase portrait and transition map of 2D system with separation line $\{p=0\}$. The parameters of numerical simulation are $G_{+}=0.3$, $G_{-}=0.075$, and $J=1.0628\times 10^{-3}$. The red solid circle and blue hollow circle are two foci, while the hollow upward triangles and downward triangles are the intersections of the trajectory and the separation line. (c)-(f) The oscillations of $u$ and $p$, respectively. (d), (f) correspond to the enlarged display of shadow region in (c), (e) respectively.
The periodic solutions of system (16) can be intuitively understood. Figure 2 illustrates a typical periodic solution, with parameters $G_{+}=0.3$, $G_{-}=0.075$, and $J=1.0628\times 10^{-3}$. For all simulations presented in this paper, the values of $G_{\pm}$ remain the same. Fig. 2(a) and 2(b) show its phase portrait and the corresponding transition map of the periodic steady state, respectively. In each half-plane, the trajectory is well-defined, spiraling clockwise around a focus. At a given moment, the trajectory intersects the separation line $p=0$ at the point $(u_{1},0)$, as shown in Fig. 2(b). The trajectory then enters the half-plane $\{p\leq 0\}$, continuing its clockwise spiral while asymptotically approaching the focus $(u_{+}^{*},0)$. It intersects the line $p=0$ again at $(u_{2},0)$. During this phase, the rotation radius relative to focus $(u_{+}^{*},0)$, which corresponds to the amplitude of the oscillation, decays exponentially, from a length of $(u_{1}-u_{+}^{*})$ to $(u_{+}^{*}-u_{2})$.
Subsequently, the system enters the half-plane $\{p>0\}$, switching its focus to the point $(u_{-}^{*},0)$ and continuing its spiral towards this new focus. Since $(u_{-}^{*},0)$ is located to the right of $(u_{+}^{*},0)$, the radial distance $(u_{-}^{*}-u_{2})$ is greater than $(u_{+}^{*}-u_{2})$. Thus, when the system switches subspaces, the rotation radius increases abruptly, compensating for the previous decay. Starting with $(u_{-}^{*}-u_{2})$ as the new rotation radius, the trajectory spirals clockwise around $(u_{-}^{*},0)$ until it intersects the line $p=0$ once more. For the system to maintain a periodic steady state, the trajectory must return to the point $(u_{1},0)$, completing one full cycle of motion.
Remarkably, $p=0$ is the line connecting the two foci. From a geometric perspective, it is evident that the trajectory in either half-plane rotates through an angle of $\pi$, as shown in Fig. 2(b). Specifically, the system spends $\frac{T_{\pm}}{2}$ in each half-plane, where $T_{\pm}=\frac{2\pi}{\Omega_{\pm}}\sim 2\pi$, with $\Omega_{\pm}=\sqrt{1-\frac{G_{\pm}^{2}}{4}}\sim 1$ for $G_{\pm}\ll 1$. Therefore, the overall period of the system’s complete motion remains approximately $2\pi$, which is near the reduced Rüchardt period $T_{0}=2\pi$. Having established the running time in each subspace, we now proceed to calculate the amplitude corresponding to the periodic solution as follows:
$u_{2}=u_{+}^{*}+(u_{+}^{*}-u_{1})e^{-\frac{G_{+}}{2}\cdot\frac{T_{+}}{2}},$ (17)
$u_{1}=u_{-}^{*}+(u_{-}^{*}-u_{2})e^{-\frac{G_{-}}{2}\cdot\frac{T_{-}}{2}},$
$u_{1}=\frac{u_{-}^{*}(1+e^{-\frac{G_{-}}{2}\cdot\frac{T_{-}}{2}})-u_{+}^{*}e^{-\frac{G_{-}}{2}\cdot\frac{T_{-}}{2}}(1+e^{-\frac{G_{+}}{2}\cdot\frac{T_{+}}{2}})}{1-e^{-(\frac{G_{+}}{2}\cdot\frac{T_{+}}{2}+\frac{G_{-}}{2}\cdot\frac{T_{-}}{2})}}\sim\frac{2(u_{-}^{*}-u_{+}^{*})}{\pi(G_{+}+G_{-})},\quad if\quad G_{\pm}\ll 1$ (18)
Substituting the parameters ($J,G_{\pm})$ used in Fig. 2, we calculate the following results:
$\frac{T_{+}}{2}=3.1775,\quad\frac{T_{-}}{2}=3.1438,\quad T=6.3214,$ (19)
and
$u_{1}=11.1352,\quad u_{2}=-10.7759.$ (20)
These results agree exactly with the numerical simulation shown in Fig. 2(d).
According to Eq.(18) and Eqs.(17), when $G_{\pm}\ll 1$, the rotation radius of the trajectory can be approximated as $\sim\frac{u_{1}-u_{2}}{2}\sim u_{1}-u_{+}^{*}>u_{1}\gg u_{-}^{*}-u_{+}^{*}$. Even if the separation line (i.e., $p=p_{1}\neq 0$}) is not the line determined by the two foci, the conclusion that the trajectory’s average rotation radius is significantly larger than the distance between the two foci remains valid. This indicates that, over one period, the trajectory can be approximated as a complete rotation around an equivalent point, corresponding to an angle of $2\pi$. Consequently, the system’s period remains close to the Rüchardt period, although the deviation becomes slightly more pronounced as $p_{1}\neq 0$.
For instance, Figure 3 illustrates the periodic solution when the separation line is given by $p=0.003$. Within a complete cycle, the system resides in the half-planes $p\leq 0$} and p>0{p>0} for durations of $3.9815$ and $2.4104$, respectively, with corresponding rotation angles of $1.2530~\pi$ and $0.7667~\pi$. The total rotation angle for the cycle is approximately $2.0197~\pi$, demonstrating that the trajectory still maintains a near-complete $2\pi$ rotation despite the shifted separation line.
Figure 3: (a) and (b) The phase portrait and transition map of 2D system with separation line $\{p=p_{1}\neq 0\}$. The parameters of numerical simulation are $G_{+}=0.3$, $G_{-}=0.075$, and $J=1.0628\times 10^{-3}$. The red solid circle and blue hollow circle are two foci, while the hollow upward triangles and downward triangles are the intersections of the trajectory and the separation line. (c) The oscillations of $u$. (d) The enlarged display of shadow region in (c).
IV.3 The original 3D PWLD system with one separation plane y=0
Returning to the original system described by Eqs. 10, we now consider the case in 3D space with a separation plane $\{y=0\}$. For the most typical and simplest periodic solution, the trajectory crosses the plane $\{y=0\}$ once upwards and once downwards within one period, intersecting it at two distinct points, denoted as $A$ and $B$. On the projection plane $(u,p)$, let $A^{*}$ and $B^{*}$ represent the projections of points $A$ and $B$, respectively. At these two points, the projection of the periodic orbit transitions between subspaces. This correspondence allows the separation plane $\{y=0\}$ in the 3D system to be effectively reduced to a straight line connecting $A^{*}$ and $B^{*}$ in the 2D system. As a result, the periodic solutions of the original 3D system can be studied entirely within this simplified 2D framework. Since the periodic behavior of such 2D PWLD systems has been analyzed in the previous section, the periodic properties of the 3D system can be deduced accordingly. Specifically, a periodic solution that alternates between the half-spaces $y\geq 0$} and y<0{y<0} during a complete cycle retains a period that closely approximates the Rüchardt period.
Figure 4: (a) and (b) The phase portrait and transition map of 3D system with separation plane $\{y=0\}$. The parameters of numerical simulation are $G_{+}=0.3$, $G_{-}=0.075$, and $J=1.0628\times 10^{-3}$. Under the setting of these parameters, the oscillation of displacement $y$ is symmetrical with respect to plane $\{y=0\}$. (c)-(h) The oscillations of $y$, $u$ and $p$, respectively. (d), (f) and (h) correspond to the enlarged display of shadow region in (c), (e) and (g) respectively.
Figure 5: The phase portrait, transition map and $y$, $u$ oscillation of 3D system with separation plane $\{y=0\}$. The simulation parameters corresponding to the figures in the left column are $G_{+}=0.3$, $G_{-}=0.075$, and $J=1.2\times 10^{-3}$. The simulation parameters corresponding to the figures in the right column are $G_{+}=0.3$, $G_{-}=0.075$, and $J=0.9\times 10^{-3}$. under the setting of these parameters, the oscillation of displacement $y$ is asymmetrical with respect to plane $\{y=0\}$.
Figures 4 and 5 present the numerical simulation results of the 3D system (10) under different $J$ (gas pumping). Figures 4 shows the case of symmetric oscillation of the displacement $y$ of the small ball relative to the position of the hole. When the ball passes through the small hole position, the pressure inside the container approaches the equilibrium pressure $P_{0}$, corresponding to the dimensionless reduced pressure $p\sim 0$. As seen in Fig. 4(b), in the projection plane, the equivalent separation line nearly coincides with the line connecting the two foci, corresponding to the motion depicted in Fig. 2. Over a complete period, the trajectory rotates through angles of $0.9921~\pi$ and $0.9996~\pi$ in the two half-planes, respectively. The corresponding running time 3.1524 and 3.1424 are very close to $\frac{T_{+}}{2}~(3.1778)$ and $\frac{T_{-}}{2}~(3.1438)$, respectively, and the entire period $T=6.2948$ is very close to the reduced Rüchardt period of $2\pi$. Additionally, the values of $u_{1}$ (11.1352) and $u_{2}$ (-10.7759) calculated by Eq.(20) are also very close to those in Fig. 4 (f) (11.1144 and -10.7532).
Figures 5 illustrates two cases of asymmetric oscillation of the displacement $y$. In these cases, the equivalent separation line deviates from the line connecting the two foci, corresponding to the motion depicted in Fig. 3. Over a complete period, the trajectory no longer rotates through angles close to $\pi$ in each half-plane. However, the total angular displacement still approaches $2\pi$, and the entire period (6.305 or 6.2841) remains close to the reduced Rüchardt period of $2\pi$.
In Koehler’s experiment, it is essential to adjust the gas flow rate to ensure that the small ball oscillates symmetrically about the position of the hole. In the theoretical analysis presented in Koehler’s paper Koehler1950 , it is assumed that the small ball spends equal amounts of time above and below the hole during each oscillation cycle. This assumption is necessary for carrying out the integration required for Fourier series analysis. The analysis in our paper offers a clearer understanding of the rationale behind this requirement, avoiding the need for complex mathematical calculations.
V Conclusion
In this paper, we revisit Koehler’s experiment and provide a concise theoretical interpretation of the self-sustained oscillations observed in the system. Without resorting to complex calculations, our analysis demonstrates that the oscillation frequency in Koehler’s experiment approximates the Rüchardt frequency. After explicitly presenting the governing equations of the model, we found that it is challenging to mathematically determine the precise parameter conditions under which the system exhibits periodic solutions close to the Rüchardt frequency. However, we prove that if such periodic solutions exist, specifically, solutions in which the small ball passes through the hole position once upward and once downward during a single period, then the system’s motion frequency is indeed very close to the Rüchardt frequency.
The self-sustained oscillations in Koehler’s experiment differ fundamentally from simple harmonic or linearly forced oscillations, representing an early nonlinear phenomenon typically encountered by undergraduates. While students can complete the experimental procedures without distinguishing between the principles underlying Koehler’s and the Rüchardt experiments, we believe that for students proficient in observation and inquiry, exploring the theoretical issues behind Koehler’s experiment can deepen their understanding of periodic motion and offer a practical scenario for studying dynamical systems.
Author Declarations
Conflict of Interest
The authors have no conflicts of interest to disclose.
Appendix A Details of our experimental setup
Figure A1 shows the details of our experimental setup. The key parameters are as follows: The volume of vessel A is $V_{0}=2.5\,\mathrm{L}$, The ball is a stainless steel ball with a diameter of 14.00 mm and a mass of 11.40 g. The inner diameter of the glass tube is 14.05 mm. The outer ends of the glass tube and the fine gas inlet tube are connected to vessel A using frosted glass surfaces to ensure airtightness while facilitating assembly and disassembly. The hole in the wall of the tube was drilled with a wide-angle drill, with an inner diameter of approximately 1.0 mm and an outer diameter of approximately 2.5 mm. As illustrated in the figure, the specific shape of the hole is not a necessary requirement but is merely intended to facilitate manufacturing. If adjustment of the hole size is required, a thin sheet with a smaller aperture can be used to block the original hole. The air pump we employed is an aquarium air pump, model SB-748, manufactured by Zhongshan SOBO Electric Appliance Co., Ltd. (China), with a power rating of 8 W and a maximum air output of 7 L/min. Vessel B serves as a gas reservoir, functioning to buffer and reduce pressure, thereby mitigating the effects of uneven air delivery from the air pump.
Figure A1: Details of our experimental setup.
Appendix B Estimating the values of parameters in Eqs. (9)
Figure (A2) shows the variation of the air pressure inside the container over time during the self-sustained oscillations of the ball, measured by a miniature wireless Bluetooth barometer (SPL06-001 pressure sensor, Goertek Microelectronics Inc., China) placed inside the vessel A. The atmospheric pressure $B=92400\,\mathrm{Pa}$. It can be seen that the equilibrium pressure $P_{0}\sim 93155\,\mathrm{Pa}$, the period $T\sim 0.6\,\mathrm{s}$ and the angular frequency $\omega\sim 10\,\mathrm{rad/s}$. The dimensionless parameter $p_{0}=\frac{P_{0}}{B}=1.008$, $A=\frac{\pi RB}{m\omega^{2}}=1782$. The fluctuation of pressure within half a period $T/2$ is approximately $170\,\mathrm{Pa}$, so $\frac{\partial P}{\partial t}=jB-g(x)(P-B)$ is of the order of magnitude of $\frac{170\,\mathrm{Pa}}{0.3\,\mathrm{s}}$. Considering that the contributions from gas supply and leakage are comparable in scale, it is appropriate to assign gas leakage parameters
$e(P_{0}-B)\sim\frac{170\,\mathrm{Pa}}{0.3\,\mathrm{s}},$
$(e+f)(P_{0}-B)\sim 3\times\frac{170\,\mathrm{Pa}}{0.3\,\mathrm{s}}.$
Then we obtain
$e\sim 0.75\,\mathrm{s}^{-1},\,f\sim 1.5\,\mathrm{s}^{-1},$
the dimensionless
$G(y)=\begin{cases}G_{+}=(e+f)/\omega=0.25,\quad&y\geq 0,\\ G_{-}=e/\omega=0.075,\quad&y<0.\end{cases}$
As provided in our analysis section, for fixed point $(u^{*},p^{*})$,
$u^{*}=A\left[J-G_{\pm}(p_{0}-1)\right].$
Let $J$ (gas supply parameter) as the control parameter, to generate self-sustained oscillations, the value of $J$ must be such that
$u^{*}_{+}=A\left[J-G_{+}(p_{0}-1)\right]<0$ (the ball’s motion tends towards the steady state of falling),
$u^{*}_{-}=A\left[J-G_{-}(p_{0}-1)\right]>0$ (the ball’s motion tends towards the steady state of rising).
Figure A2: The variation of the air pressure inside the container over time during the self-sustained oscillations of the ball, in an actual experiment.